OpenStudy (anonymous):

Given that g(x) = 17x - x2; find all values of a such that g(a) = 72

OpenStudy (anonymous):

g(a)=72=17a-a^2 -a^2 + 17a -72=0 a^2 -17a +72=0 solve quadratic equation for a

OpenStudy (sriram):

g(a)=17a-a^2=72 a^2-17a+72=0 a=9 or a=8